# Breaking Down the Barriers: Investigating Non-Expert User Experiences in Robotic Teleoperation in UK and Japan

### v1

@@ Abstract
Robots are being created each year with the goal of integrating them into our daily lives. As such, there is an interest in research in evaluating the trust of humans toward robots. In addition, teleoperating robotic arms can be challenging for non-experts. In order to reduce the strain put on the user, we created TELESIM, a modular and plug-and-play framework that enables direct teleoperation of any robotic arm using a digital twin as the interface between users and the robotic system. However, analysis of the strain put on the user and its ability to trust robots was omitted. This paper addresses these omissions by presenting the additional results of our user survey of 37 participants carried out in UK. In addition, we present the results of an additional user survey, under similar conditions performed in Japan, with the goal of addressing the limitations of our previous approach, by interfacing a VR controller with a UR5e. Our experimental results show that the UR5e has a higher number of towers built. Additionally, the UR5e gives the least amount of cognitive stress, while the combination of Senseglove and UR3 gives the user the highest physical strain and causes the user to feel more frustrated. Finally, Japanese seems more trusting towards robots than British.
@@ I Introduction
The advent of Industry 4.0 has fundamentally transformed the manufacturing landscape which has marked a transition from conventional programmable robots to sophisticated, data-driven systems [1, 2]. This paradigm shift, coupled with advancements in digital twin technology, has enabled the development of virtual replicas of smart factories in order to enhance operational efficiency and decision-making processes. As we progress into the era of Industry 5.0, the focus is now on fostering an ecologically and socially responsible industry that prioritizes human-centric values [3]. This evolution needs the emergence of machines that extend beyond mere digital replicas. That is, these machines have to evolve into collaborative partners that support humans in navigating complex tasks.
Central to the framework of Industry 5.0 are teleoperation systems, which facilitate remote human-machine collaboration [3]. However, extended teleoperation sessions present challenges, notably leading to physical and mental fatigue among operators[4]. Moreover, the requisite level of expertise for effective teleoperation remains inadequately defined[5]. This highlights the need for further research to establish standardized protocols and training programs to enhance teleoperation practitioners’ efficiency and well-being. Therefore, This paper investigates the influence of the robotic hardware and participant demographics on performance metrics and user workload in the context of direct teleoperation.
Our previous work introduced TELESIM, a framework for intuitive robotic arm teleoperation[5]. While our initial study with over 30 participants demonstrated that non-experts could effectively teleoperate robotic arms, it did not fully address the physical and cognitive strain experienced by users. This critical aspect warrants further investigation to ensure user safety and well-being in teleoperation scenarios[4].
@@
To address these gaps and expand on our previous findings, this paper rigorously explores teleoperation’s impact on users’ mental and physical health and their relationship with robots during task performance. For this, we conducted a large-scale, international user survey across Japan and the UK, involving over 70 participants from diverse backgrounds. The study was carried out using three robots with varying ranges and speeds: the Universal Robot 3, Universal Robot 5e, and Rethink Robotics Baxter. Participants performed a standardized 3-cube tower stacking task for 10 minutes which allowed us to direct comparison across robot types and user demographics. Our contributions are thus the following:
@@
    A large-scale, cross-cultural comparison of teleoperation performance and user experience between Japan and the UK, involving 74 participants from diverse backgrounds.
@@
    An in-depth analysis of the relationship between robot capabilities (e.g., reach, stability, control method) and teleoperation performance, derived from experiments with three different robotic arms: Baxter, UR3, and UR5e.
@@
    A comprehensive evaluation of user workload during teleoperation using the NASA-TLX questionnaire. We found that the difference in hardware has an impact on the mental workload and that the user’s frustration is not linked to the performance of the teleoperation.
@@
    An investigation into the differences in trust towards robots between Japanese and British participants, challenging previous findings that British users trusted robots more than Japanese[6].
The rest of this paper is organised as follows. In Section II, we explore the current state of the art and the research gap. Specifically, in Section II-A, we review existing teleoperation frameworks and the methodologies other researchers have employed to evaluate them. Then, in Section II-B, we explore ways the user’s workload during teleoperation has been assessed. Finally, in Section II-C, we inspect how trust has been recorded and evaluated in teleoperation. Afterwards, in Section III we present a recapitulation of the TELESIM framework and our experimental design implemented in the United Kingdom. Thus, we detail in Section IV a detailed description of our experimental setup, with particular emphasis on the distinctions between our experiments conducted in Japan and the United Kingdom. Finally, in Section V, we provide a comprehensive analysis and comparison of the workload and performance of our teleoperation framework.
@@ II-A Teleoperation Systems
Direct teleoperation is recognized as an essential precursor to shared autonomy [7]. This recognition arises from the considerable cognitive demands placed on users during direct teleoperation [4] and the challenges users face in executing precise, millimeter-scale adjustments to a robot’s end effector. In medical applications, user movements are often scaled down to enhance precision [8, 9]. However, such scaling is not universally applicable across all manipulation tasks, particularly those requiring extensive arm movements for object relocation.
While there has been substantial research in the field, there are no studies that evaluate direct teleoperation across multiple robots without incorporating shared autonomy. Existing literature primarily focuses on shared autonomy to alleviate cognitive strain while maximizing precision. For example, researchers have explored various control methodologies, such as low degree-of-freedom interfaces such as keyboards [10], joysticks [11, 12], touchscreens [13], and gamepads [14], which have shown improved control levels and reduced mental strain [10]. With the advent of virtual reality (VR) and augmented reality (AR) technologies, VR controllers [15, 16] and smartphones [17] have been investigated for their applicability in teleoperation. Some studies have also examined motion mapping of the user’s body [18, 19] or gaze control [20], though these methods present challenges such as kinematic discrepancies and increased cognitive load [4]. While extensive research has been conducted, the diversity of findings allows for tailored control methods that can be selected based on specific cases. The control method, therefore, remains a choice made on a case-by-case basis.
To address the limitations of direct teleoperation, researchers have emphasized the benefits of shared autonomy in enhancing task success rates. However, the absence of literature focusing solely on direct teleoperation suggests that these limitations may be attributed to the specific implementation of teleoperation in each study. Notably, researchers have employed either MoveIt! [21], which has its own limitations, described in Section III, or custom planning interfaces for robot control, which is either not detailed how teleoperation was implemented [22], or whether teleoperation was developed for the specific shared autonomy framework[23, 24]. In addition, these studies have predominantly focused on single robotic systems and conducted experiments with small user groups (between 8 and 23 participants), emphasizing different autonomy levels rather than control methods.
@@ II-B Evaluating Workload for Teleoperation Systems
Shared autonomy has the potential to enhance the performance of direct teleoperation, but human factors significantly influence the overall success of tasks. Therefore, an efficient teleoperation system must minimize both physical and cognitive stress on the operator, and assessing user workload is crucial, with the NASA Task Load Index (NASA-TLX) [25] being the predominant tool utilized in the literature for this evaluation. The NASA-TLX questionnaire was developed by Sandra Hart at the NASA Ames Research Center in 1988[26], is a widely used subjective assessment tool specifically designed to measure perceived workload across various tasks and environments. It evaluates workload based on six distinct dimensions: Mental Demand, Physical Demand, Temporal Demand, Performance, Effort, and Frustration. Two versions of the NASA-TLX are available. The original version includes six 21-point scales for each dimension and a pairwise comparison of each dimension, whereas the ”Raw” version omits the pairwise comparison [25]. The ”Raw” version is often favored because it is easier to carry out and analyze[25].
In robotic teleoperation, the NASA-TLX is the preferred instrument for assessing workload[22, 27, 24]. To the best of our knowledge, only a few research works have opted not to use the NASA or ”Raw” NASA questionnaire for robotic teleoperation. That is, some researchers instead decided to use the NASA questionnaire as a way to measure the user’s workload using extra devices or metrics. For example, Moya et al. [28] utilized EEG but corroborated these results with the NASA questionnaire. Additionally, researchers such as Naughton et al. [29] were interested solely in the overall NASA score, which they obtained by summing all dimensions. However, it is not the only way to use the NASA questionnaire. Parsa et al. [27] conducted an ANOVA analysis on each NASA dimension rather than using the questionnaire score to extract more value out of the questionnaire. Others, such as Lin et al. [22], focused on specific metrics to emphasize particular results[30]. Finally, Bechtel et al.[31] employed a single question for users to self-report their overall workload on a Likert-Scale, similar to the Single Ease Question, which is ”Overall, This task was ?”
@@ II-C Trust in Robots
Trust in robots is a critical factor, especially given the recent increase in robots designed to interact with the general population[32]. Consequently, researchers have developed various questionnaires tailored to different use cases, as evaluating trust is complex and often requires multiple scales[33]. Prominent scales include the General Attitudes Towards Robots Scale (GAToRS)[34], the Robotic Social Attributes Scale (RoSAS)[35], and the Negative Attitude Towards Robots Scale (NARS)[36]. These scales are designed to evaluate social robots with which users can interact physically or verbally, or establish an anthropomorphic connection. NARS [36] remains popular due to its three subscales, which can analyze a broad range of attitudes toward robots: negative attitudes toward interaction situations with robots (S1), negative attitudes toward the social influence of robots (S2), and negative attitudes toward emotions in interaction with robots (S3). Each scale can be employed individually to examine specific cases. Notably, NARS is the only trust scale used to evaluate trust in robotic arm manipulation[37].
@@ III Teleoperation Framework
TELESIM [5] (Fig. 2) is a framework that enables teleoperation of various robotic arms. The framework comprises three components, with the controller (1) being any control system that outputs a 3D pose (described in Sec. III-A. TELESIM, (2) in Figure 2 and described in Sec. III-B, computes a real-time motion plan for the pose provided by the controller (1) using Nvidia Isaac Sim [38], an RTX simulator serving as a digital twin of the environment. The joint states of the robot are then sent through ROS2 to any robotic arm that supports ROS2 control, (3) in Fig. 2.
@@ III-A Controllers
Our choice of control method was informed by the work of Gottardi et al. [24] who explored integrating multiple control systems, including the combination of VR controllers with tracking bands to monitor user movements and Rakita et al. [9] who compared several control methods and integrated them into a custom inverse kinematics solver to align the end-effector pose with the user’s input. Their findings indicated a preference for VR controllers in completing pick-and-place tasks. Thus, the controller for TELESIM, (2) in Fig. 2, consisted of a Steam Index VR controller used to control a Rethink Robotics Baxter robot and the UR5e and a Senseglove to control the UR3. We used the Senseglove for mapping individual finger motions, particularly the index and thumb, to the two fingers of the Yale T42 gripper, described in Section III-C. A Vive VR Tracker, mounted on top of the Senseglove, was employed to map user motions into the UR3 robot.
The Senseglove functions analogously to a glove affixed to the palm and fingertips. For our user survey, we opted for the Senseglove development kit (shown in Figure 2(1)) due to its ease of application and adjustment. Moreover, we limited the strapping to the index finger and thumb, corresponding to our two-fingered gripper design. Although the Senseglove SDK is compatible with multiple platforms, integration into our framework required us to port it to ROS2. This was accomplished through the creation of a ROS2 control plugin [^1] https://github.com/09ubberboy90/senseglove_ros2_ws. This adaptation enabled the seamless incorporation of the Senseglove functionality within our TELESIM framework. Using this plugin, we are able to read the state of the user’s hand by aggregating the joint angles of each fingers and scaling them to a range corresponding to the gripper finger being fully opened and fully closed. This scaling aligns with the motor control range, derived from average values recorded in practice for a fully opened hand and a closed hand. This methodology enables a standardized mapping of finger positions to motor control inputs, facilitating the accurate translation of user hand movements to gripper actions.
@@ III-B TELESIM
The robotic systems are dynamically instantiated in Isaac Sim from their Universal Robot Descriptor Files (URDF) to enhance flexibility and adaptability to any robot and make TELESIM a modular and plug-and-play framework. Isaac Sim offers a diverse array of methodologies for motion planning and collision avoidance. For instance, Rapidly Exploring Random Tree (RRT) algorithms[41], specifically Nvidia’s Lula RRT implementation, offer global solutions primarily suited for static environments. Conversely, Moveit2[42], the principal motion planning library for ROS2, utilizes the Open Motion Planning Library (OMPL)[43]. While OMPL yields global solutions, its efficacy in real-time and dynamic applications, executed through MoveIt Servo [^2] https://moveit.picknik.ai/main/doc/examples/realtime_servo/-realtime_servo_tutorial.html, frequently results in singularities, necessitating robot movement away from objectives for recovery. We instead chose Riemannian Motion Policy (RMP)[44] which is distinctively engineered for dynamic environments and supports real-time operation and collision avoidance.
@@ III-C Robots and End-effectors
For our experiment we used 3 different grippers, Baxter robot default linear gripper, depicted in Figure 1, used with the Baxter robot, a modified Yale T42 gripper[45], shown in Figure 3, used with the UR3 robot and a Robotiq 2F-140 gripper [46], shown in Figure 5, used with the UR5e robot. The difference in gripper was informed due to our belief that hardware played a part in teleoperation performance. These 3 grippers have varying heights and opening lengths. The Baxter gripper was chosen as a pure linear gripper that is fully integrated with the Baxter robot. The Yale T42 gripper is a 2-finger gripper designed with individual finger motion, causing a significant difference in height between the open and close state. Finally, the Robotiq is a blend of a liner gripper and a finger gripper, giving it a larger opening between the 2 fingers at the cost of a change in height between its 2 states.
@@
The Baxter gripper is controlled using the same control method as Baxter, which is a ROS2 control package, communicating with the real robot using a ROS2 to ROS1 bidirectional bridge using the joints states provided by Isaac Sim following the setup described in Section III-B. This package was adapted from the ROS1 package provided by Rethink Robotics to ROS2333https://github.com/CentraleNantesRobotics/baxter_common_ros2 and modified to allow more control of the gripper. The change of gripper state is controlled by a custom ROS2 node listening to the state of the VR controller, as described in Section III-A and publishes a trigger to the digital clone in Isaac Sim whenever the trigger button of the controller is pressed. Thus, this causes the digital clone joint states to be updated and forwarded to the real robot using the ROS2 control package.
@@
The Yale T42 gripper was also chosen due to its 3D printable components and individually controllable fingers, each utilizing two Dynamixel motors. This configuration aligned with the requirement of mapping individual finger movements to the Senseglove interface. An interface was developed to address the lack of ROS2 integration in the SDK. This involved developing a ROS2 package that incorporates the SDK into a ROS2 control interface to enable seamless integration with TELESIM.
@@
Initial testing revealed limitations in the control board’s communication capabilities. The single high-speed UART channel resulted in significant latency between issuing a command and its execution. Furthermore, the intended implementation of haptic feedback from motor data to the Senseglove required a higher data transmission rate than the virtual UART communication could support. To address these issues, a dual Arduino board setup employing I2C communication was implemented, effectively resolving the bandwidth constraints. The decision was made to forgo haptic feedback implementation for this paper, as it would have been the sole control method featuring this capability, potentially complicating comparisons between control methods in user surveys and experiments. Although vibration on the Vive controller could theoretically serve as a form of haptic feedback, this approach would introduce additional variables requiring evaluation, such as vibration strength and delay, to assess their impact on user performance and perception.
@@
By incorporating two Arduino boards in the gripper, we had to alter the original gripper’s mechanical design. A secondary level was introduced between the motor and the robot attachment point to accommodate the control boards. This change is shown in Fig. 3. This solution was adopted after attempts to mount the boards on the gripper’s side which resulted in cable entanglement and damage during robot operation due to spatial constraints. While the additional level resolved the cable management issues, it extended the gripper’s length by 5 cm, consequently reducing the UR3 robot’s already limited workspace.
@@
The finger control mechanism uses a fishing line under tension, driven by the motor. Initial tests revealed that the original fishing line lacked the necessary strength for teleoperation tasks, frequently breaking during use. Although stronger fishing lines were considered, they imposed excessive strain on the motors. To preserve motor integrity, the decision was made to retain the weaker fishing line despite its limitations. The final modification involved the removal of the final degree of freedom on each finger. This alteration was implemented as the original finger configuration hindered the grasping of cubic objects without providing significant advantages for the objectives of this paper. The modifications to the Yale T42 gripper, while addressing critical functional requirements, introduced trade-offs in terms of workspace reduction and grasping force limitations. These adaptations were necessary to meet the specific demands of the teleoperation experiment within the constraints of the available hardware used in our experiments in Sec. IV.
It is controlled using a custom ROS2 package, that listen to the joint states given by the digital clone in Isaac Sim and send them through I2C to the first Arduino board and republishes the states of the motor, if needed to provide an optional force feedback mechanism to the Senseglove. This packages runs in addition to a ROS2 control plugin provided by the Universal Robot company [^4] https://github.com/UniversalRobots/Universal_Robots_ROS2_Driver, needed to control the both the UR3 and the UR5e.

### v2

@@ Abstract
Robots are being created each year with the goal of integrating them into our daily lives. As such, there is an interest in research in evaluating the trust of humans toward robots. In addition, teleoperating robotic arms can be challenging for non-experts. To reduce the strain put on the user, we created TELESIM, a modular and plug-and-play framework that enables direct teleoperation of any robotic arm using a digital twin as the interface between users and the robotic system. We evaluated our framework using a user survey with three robots and control methods and recorded the user’s workload and performance at completing a tower stacking task. However, an analysis of the strain on the user and their ability to trust robots was omitted. This paper addresses these omissions by presenting the additional results of our user survey of 37 participants carried out in United Kingdom. In addition, we present the results of an additional user survey, under similar conditions performed in Japan, with the goal of addressing the limitations of our previous approach, by interfacing a VR controller with a UR5e. Our experimental results show that the UR5e has more towers built. Additionally, the UR5e gives the least amount of cognitive stress, while the combination of Senseglove and UR3 provides the user with the highest physical strain and causes the user to feel more frustrated. Finally, the Japanese participants seem more trusting of robots than the British participants.
@@ I Introduction
The advent of Industry 4.0 has fundamentally transformed the manufacturing landscape which has marked a transition from conventional programmable robots to sophisticated, data-driven systems [1, 2]. This paradigm shift, coupled with advancements in digital twin technology, has enabled the development of virtual replicas of smart factories to enhance operational efficiency and decision-making processes. As we progress into the era of Industry 5.0, the focus is now on fostering an ecologically and socially responsible industry that prioritizes human-centric values [3]. This evolution needs the emergence of machines that extend beyond mere digital replicas. That is, these machines have to evolve into collaborative partners that support humans in navigating complex tasks.
Central to the framework of Industry 5.0 is teleoperation systems, which facilitate remote human-machine collaboration [3]. However, extended teleoperation sessions present challenges, notably leading to physical and mental fatigue among operators [4]. Moreover, the requisite level of expertise for effective teleoperation still needs to be more adequately defined [5]. This highlights the need for further research to establish standardized protocols and training programs to enhance teleoperation practitioners’ efficiency and well-being. Therefore, This paper investigates the influence of robotic hardware and participant demographics on performance metrics and user workload in the context of direct teleoperation.
Our previous work introduced TELESIM, a framework for intuitive robotic arm teleoperation [5]. While our initial study with over 30 participants demonstrated that non-experts could effectively teleoperate robotic arms, it did not fully address the physical and cognitive strains experienced by users. This critical aspect warrants further investigation to ensure user safety and well-being in teleoperation scenarios [4].
@@
To address these gaps and expand on our previous findings, this paper rigorously explores teleoperation’s impact on users’ mental and physical health and their relationship with robots during task performance. For this, we conducted a large-scale, international user survey across Japan and the United Kingdom, involving over 70 participants from diverse backgrounds. The study used three robots with varying ranges and speeds: the Universal Robot 3, Universal Robot 5e, and Rethink Robotics Baxter. Participants performed a standardized 3-cube tower stacking task for 10 minutes, which allowed us to compare robot types and user demographics directly. Our contributions are thus the following:
@@
    A large-scale, cross-cultural comparison of teleoperation performance and user experience between Japan and the United Kingdom (UK), involving 74 participants from diverse backgrounds.
@@
    An in-depth analysis of the relationship between robot capabilities (e.g., reach, stability, control method) and teleoperation performance derived from experiments with three robotic arms: Baxter, UR3, and UR5e.
@@
    A comprehensive evaluation of user workload during teleoperation using the NASA-TLX questionnaire. We found that the difference in hardware impacts the mental workload and that the user’s frustration is not linked to the performance of the teleoperation.
@@
    An investigation into the differences in trust towards robots between the Japanese and the British participants, challenging previous findings that the British users trusted robots more than the Japanese participants [6].
The rest of this paper is organized as follows. In Section II, we explore the current state of the art and the research gap. Specifically, in Section II-A, we review existing teleoperation frameworks and the methodologies other researchers have employed to evaluate them. Then, in Section II-B, we explore ways the user’s workload during teleoperation has been assessed. Finally, in Section II-C, we inspect how trust has been recorded and evaluated in teleoperation. Afterwards, in Section III, we present a recapitulation of the TELESIM framework and our experimental design implemented in the United Kingdom. Thus, we detail in Section IV a detailed description of our experimental setup, emphasizing the distinctions between our experiments conducted in Japan and the United Kingdom. Finally, in Section V, we provide a comprehensive analysis and comparison of the workload and performance of our teleoperation framework.
@@ II-A Teleoperation Systems
Direct teleoperation is recognized as an essential precursor to shared autonomy [7]. This recognition arises from the considerable cognitive demands placed on users during direct teleoperation [4] and the challenges users face in executing precise, millimetre-scale adjustments to a robot’s end effector. In medical applications, user movements are often scaled down to enhance precision [8, 9]. However, such scaling is only universally applicable across some manipulation tasks, particularly those requiring extensive arm movements for object relocation.
While substantial research has been done in the field, only our previous work [5] has evaluated direct teleoperation across multiple robots without incorporating shared autonomy. Existing literature primarily focuses on shared autonomy to alleviate cognitive strain while maximizing precision. For example, researchers have explored various control methodologies, such as low degree-of-freedom interfaces such as keyboards [10], joysticks [11, 12], touchscreens [13], and gamepads [14], which have shown improved control levels and reduced mental strain [10]. With the advent of virtual reality (VR) and augmented reality (AR) technologies, VR controllers [15, 16] and smartphones [17] have been investigated for their applicability in teleoperation. Some studies have also examined motion mapping of the user’s body [18, 19] or gaze control [20], though these methods present challenges such as kinematic discrepancies and increased cognitive load [4]. While extensive research has been conducted, the diversity of findings allows for tailored control methods that can be selected based on specific cases. The control method, therefore, remains a choice made on a case-by-case basis.
To address the limitations of direct teleoperation, researchers have emphasized the benefits of shared autonomy in enhancing task success rates. However, the absence of literature focusing solely on direct teleoperation suggests that these limitations may be attributed to the specific implementation of teleoperation in each study. Notably, researchers have employed either MoveIt!  [21], which has its limitations, described in Section III, or custom planning interfaces for robot control, which is either not detailed how teleoperation was implemented [22], or whether teleoperation was developed for the specific shared autonomy framework[23, 24]. In addition, these studies have predominantly focused on single robotic systems and conducted experiments with small user groups (between 8 and 23 participants), emphasizing different autonomy levels rather than control methods.
@@ II-B Evaluating Workload for Teleoperation Systems
Shared autonomy can potentially enhance direct teleoperation performance, but human factors significantly influence the overall success of tasks. Therefore, an efficient teleoperation system must minimize both physical and cognitive stress on the operator, and assessing user workload is crucial, with the NASA Task Load Index (NASA-TLX) [25] being the predominant tool utilized in the literature for this evaluation. Sandra Hart developed the NASA-TLX questionnaire at the NASA Ames Research Center in 1988 [26], a widely used subjective assessment tool designed to measure perceived workload across various tasks and environments. It evaluates workload based on six dimensions: Mental Demand, Physical Demand, Temporal Demand, Performance, Effort, and Frustration. Two versions of the NASA-TLX are available. The original version includes six 21-point scales for each dimension and a pairwise comparison of each dimension, whereas the ”Raw” version omits the pairwise comparison [25]. The ”Raw” version is often favoured because it is easier to carry out and analyse [25].
In robotic teleoperation, the NASA-TLX is the preferred instrument for assessing workload [22, 27, 24]. To our knowledge, only a few research works have opted not to use the NASA or ”Raw” NASA questionnaire for robotic teleoperation. That is, some researchers instead decided to use the NASA questionnaire to measure the user’s workload using extra devices or metrics. For example, Moya et al. [28] utilized EEG but corroborated these results with the NASA questionnaire. Additionally, researchers such as Naughton et al. [29] were interested solely in the overall NASA score, which they obtained by summing all dimensions. However, it is not the only way to use the NASA questionnaire. Parsa et al. [27] conducted an ANOVA analysis on each NASA dimension rather than using the questionnaire score to extract more value from the questionnaire. Others, such as Lin et al. [22], focused on specific metrics to emphasize particular results [30]. Finally, Bechtel et al. [31] employed a single question for users to self-report their overall workload on a Likert-Scale, similar to the Single Ease Question, which is ”Overall, This task was ?”
@@ II-C Trust in Robots
Trust in robots is a critical factor, especially given the recent increase in robots designed to interact with the general population [32]. Consequently, researchers have developed various questionnaires tailored to different use cases, as evaluating trust is complex and often requires multiple scales [33]. Prominent scales include the General Attitudes Towards Robots Scale (GAToRS) [34], the Robotic Social Attributes Scale (RoSAS) [35], and the Negative Attitude Towards Robots Scale (NARS) [36]. These scales evaluate social robots with which users can interact physically or verbally or establish an anthropomorphic connection. NARS [36] remains popular due to its three subscales, which can analyse a broad range of attitudes toward robots: negative attitudes toward interaction situations with robots (S1), negative attitudes toward the social influence of robots (S2), and negative attitudes toward emotions in interaction with robots (S3). Each scale can be employed individually to examine specific cases. NARS is the only trust scale used to evaluate trust in robotic arm manipulation [37].
@@ III Teleoperation Framework
TELESIM  [5] (Fig. 2) is a framework that enables teleoperation of various robotic arms. The framework comprises three components, with the controller (1) being any control system that outputs a 3D pose (described in Sec. III-A. TELESIM, (2) in Figure 2 and described in Sec. III-B, computes a real-time motion plan for the pose provided by the controller (1) using Nvidia Isaac Sim  [38], an RTX simulator serving as a digital twin of the environment. The joint states of the robot are then sent through ROS2 to any robotic arm that supports ROS2 control (3) in Fig. 2.
@@ III-A Controllers
Our choice of control method was informed by the work of Gottardi et al. [24], who explored integrating multiple control systems, including the combination of VR controllers with tracking bands to monitor user movements, and Rakita et al. [9] who compared several control methods and integrated them into a custom inverse kinematics solver to align the end-effector pose with the user’s input. Their findings indicated a preference for VR controllers in completing pick-and-place tasks. Thus, the controller for TELESIM, (2) in Fig. 2, consisted of a Steam Index VR controller used to control a Rethink Robotics Baxter robot and the UR5e and a Senseglove to control the UR3. We used the Senseglove for mapping individual finger motions, particularly the index and thumb, to the two fingers of the Yale T42 gripper, described in Section III-C. A Vive VR Tracker, mounted on top of the Senseglove, was employed to map user motions into the UR3 robot.
The Senseglove functions analogously to a glove affixed to the palm and fingertips. For our user survey, we opted for the Senseglove development kit (shown in Figure 2(1)) due to its ease of application and adjustment. Moreover, we limited the strapping to the index finger and thumb, corresponding to our two-fingered gripper design. Although the Senseglove SDK is compatible with multiple platforms, integration into our framework required us to port it to ROS2. This was accomplished through the creation of a ROS2 control plugin [^1] https://github.com/09ubberboy90/senseglove_ros2_ws. This adaptation enabled the seamless incorporation of the Senseglove functionality within our TELESIM framework. Using this plugin, we are able to read the state of the user’s hand by aggregating the joint angles of each finger and scaling them to a range corresponding to the gripper finger being fully opened and fully closed. This scaling aligns with the motor control range, derived from average values recorded in practice for a fully opened hand and a closed hand. This methodology enables a standardized mapping of finger positions to motor control inputs, facilitating the accurate translation of user hand movements to gripper actions.
@@ III-B TELESIM
The robotic systems are dynamically instantiated in Isaac Sim from their Universal Robot Descriptor Files (URDF) to enhance flexibility and adaptability to any robot and make TELESIM a modular and plug-and-play framework. Isaac Sim offers a diverse array of methodologies for motion planning and collision avoidance. For instance, Rapidly Exploring Random Tree (RRT) algorithms [41], specifically Nvidia’s Lula RRT implementation, offer global solutions primarily suited for static environments. Conversely, Moveit2 [42], the principal motion planning library for ROS2, utilizes the Open Motion Planning Library (OMPL) [43]. While OMPL yields global solutions, its efficacy in real-time and dynamic applications, executed through MoveIt Servo [^2] https://moveit.picknik.ai/main/doc/examples/realtime_servo/-realtime_servo_tutorial.html, frequently results in singularities, necessitating robot movement away from objectives for recovery. We instead chose Riemannian Motion Policy (RMP) [44], which is distinctively engineered for dynamic environments and supports real-time operation and collision avoidance.
@@ III-C Robots and End-effectors
For our experiment we used 3 different grippers, Baxter robot default linear gripper, depicted in Figure 1, used with the Baxter robot, a modified Yale T42 gripper [45], shown in Figure 3, used with the UR3 robot and a Robotiq 2F-140 gripper  [46], shown in Figure 5, used with the UR5e robot. The difference in gripper was informed due to our belief that hardware played a part in teleoperation performance. These 3 grippers have varying heights and opening lengths. The Baxter gripper was chosen as a pure linear gripper that is fully integrated with the Baxter robot. The Yale T42 gripper is a 2-finger gripper designed with individual finger motion, causing a significant difference in height between the opened and closed state. Finally, the Robotiq is a blend of a liner gripper and a finger gripper, giving it a larger opening between the 2 fingers at the cost of a change in height between its 2 states.
@@
The Baxter gripper is controlled using the same control method as Baxter, which is a ROS2 control package, communicating with the real robot using a ROS2 to ROS1 bidirectional bridge using the joints states provided by Isaac Sim following the setup described in Section III-B. This package was adapted from the ROS1 package provided by Rethink Robotics to ROS2333https://github.com/CentraleNantesRobotics/baxter_common_ros2 and modified to allow more control of the gripper. The change of gripper state is controlled by a custom ROS2 node listening to the state of the VR controller, as described in Section III-A, and publishes a trigger to the digital clone in Isaac Sim whenever the trigger button of the controller is pressed. Thus, this causes the digital clone joint states to be updated and forwarded to the real robot using the ROS2 control package.
@@
The Yale T42 gripper was also chosen due to its 3D printable components and individually controllable fingers, each utilizing two Dynamixel motors. This configuration aligned with the requirement of mapping individual finger movements to the Senseglove interface. An interface was developed to address the SDK’s lack of ROS2 integration. This involved developing a ROS2 package to incorporate the SDK into a ROS2 control interface to enable seamless integration with TELESIM.
@@
Initial testing revealed limitations in the control board’s communication capabilities. The single high-speed UART channel resulted in significant latency between issuing a command and its execution. Furthermore, the intended implementation of haptic feedback from motor data to the Senseglove required a higher data transmission rate than the virtual UART communication could support. A dual Arduino board setup employing I2C communication was implemented to address these issues, effectively resolving the bandwidth constraints. The decision was made to forgo haptic feedback implementation for this paper, as it would have been the sole control method featuring this capability, potentially complicating comparisons between control methods in user surveys and experiments. Although vibration on the Vive controller could theoretically serve as a form of haptic feedback, this approach would introduce additional variables requiring evaluation, such as vibration strength and delay, to assess their impact on user performance and perception.
@@
We had to alter the original gripper’s mechanical design by incorporating two Arduino boards in the gripper. A secondary level was introduced between the motor and the robot attachment point to accommodate the control boards. This change is shown in Fig. 3. This solution was adopted after attempts to mount the boards on the gripper’s side, resulting in cable entanglement and damage during robot operation due to spatial constraints. While the additional level resolved the cable management issues, it extended the gripper’s length by 5 cm, consequently reducing the UR3 robot’s already limited workspace.
@@
The finger control mechanism uses a fishing line under tension driven by the motor. Initial tests revealed that the original fishing line lacked the necessary strength for teleoperation tasks, frequently breaking during use. Although stronger fishing lines were considered, they imposed excessive strain on the motors. To preserve motor integrity, the decision was made to retain the weaker fishing line despite its limitations. The final modification involved the removal of the final degree of freedom on each finger. This alteration was implemented as the original finger configuration hindered the grasping of cubic objects without providing significant advantages for the objectives of this paper. The modifications to the Yale T42 gripper, while addressing critical functional requirements, introduced trade-offs in terms of workspace reduction and grasping force limitations. These adaptations were necessary to meet the specific demands of the teleoperation experiment within the constraints of the available hardware used in our experiments in Sec. IV.
The gripper is controlled using a custom ROS2 package that listens to the joint states given by the digital clone in Isaac Sim and sends them through I2C to the first Arduino board. It republishes the states of the motor if needed to provide an optional force feedback mechanism to the Senseglove. This package runs in addition to a ROS2 control plugin provided by the Universal Robot company [^4] https://github.com/UniversalRobots/Universal_Robots_ROS2_Driver, needed to control both the UR3 and the UR5e.